Find answers to the following. Write and indicate how you solved them.
 Is 1/15 equal to 4 / 30

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

1 / 15, 4 / 30
4 / 30 = (2 × 2) / (15 × 2)
= 2 / 15
∴1 / 15 ≠ 4 / 30
Hence, 1 / 15 is not equal to 4 / 30

​Thanks